The Atlantic Company sells a product with a break-even point of 5,862 sales units. The variable cost is $84 per unit, and fixed costs are $199,308.

Determine the unit sales price. Round answer to nearest whole number.
$

Determine the break-even points in sales units if the company desires a target profit of $50,048. Round answer to the nearest whole number.
units
Answer #1

1.At breakeven;Contribution=Fixed cost=$199308

Contribution/unit=(199308/5862)=$34

Contribution=Sales-Variable cost

Hence unit sale price=(34+84)=$118

2.The breakeven point would not change with change in target profits

Hence breakeve value=(5862*118)=$691716.